OpenStudy (anonymous):

Find all non-negative steady states (y*, z*) of the following system of differential equations: (see equations in the comments).

OpenStudy (anonymous):

\[dy/dt=\alpha y(1-y-\gamma z)\] \[dz/dt=\beta z(1-\mu y-z)\]

OpenStudy (anonymous):

The steady states occur at the points \((y,z)\) that make \(\dfrac{dy}{dt}=\dfrac{dz}{dt}=0\). Are \(\alpha,\beta,\gamma,\mu\) constants?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

I don't know for sure though.